OpenStudy (anonymous):

estimate the follwoing question by using a power series expansion of sin x with five non-zero terms.

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi^2} [ \sin (\sqrt{x})] dx\]

OpenStudy (anonymous):

\[\Large sinx=x-\frac{x^3}{3!}+\frac{x^5}{5!}-\frac{x^7}{7!}+\frac{x^9}{9!} \]replace x by \[\sqrt{x}\]

OpenStudy (anonymous):

\[\Large sin\sqrt{x}=\sqrt{x}-\frac{\sqrt{x^3}}{3!}+\frac{\sqrt{x^5}}{5!}-\frac{\sqrt{x^7}}{7!}+\frac{\sqrt{x^9}}{9!}\]

OpenStudy (anonymous):

\[\Large \int\limits_{0}^{\pi^2}\Large (\sqrt{x}-\frac{\sqrt{x^3}}{3!}+\frac{\sqrt{x^5}}{5!}-\frac{\sqrt{x^7}}{7!}+\frac{\sqrt{x^9}}{9!} )dx\]

OpenStudy (anonymous):

the rest is yours (:

OpenStudy (anonymous):

that makes so much more sense now, thank you! i learned this information within hours so ive been stuck haha

OpenStudy (anonymous):